# NJL-Chiral Soliton and the Nucleon Equation of State at supra-saturation density: Impact of Chiral Symmetry Restoration

### v1

@@ II Theoretical Framework
To construct the in-medium modified nucleon state and quark core state, we follow an approach similar to that adopted in the NJL version of the chiral confining model (CCM), as briefly sketched below. The chiral confining–NJL model captures the essential dynamics of chiral symmetry breaking and quark confinement, providing an effective framework to describe low-energy QCD phenomena. It elucidates the response of composite nucleons to the nuclear scalar field and the interplay between confinement and chiral symmetry breaking in the in-medium evolution of the nucleon mass. However, in the present work, we restrict our analysis to the NJL-soliton framework for describing the nucleon, without introducing an explicit confining force such as that associated with a string tension. Instead, the role of confinement is effectively mimicked by mesons (with the soliton stabilized by vector mesons in the current framework) within the soliton picture. A description of the nucleon, including an explicit confining mechanism, will be addressed in future work. As detailed in Ref. [67, 68], the CCM provides a robust theoretical framework for understanding QCD phenomenology, particularly in bridging QCD with the description of nuclear and NS matter. In its NJL version, this model builds upon the NJL model, whose bosonization procedure and theoretical formulation are comprehensively presented in Ref. [63, 68]. We adopt the same approach here.
II.1 Bosonized NJL Model and Chiral Soliton Ansatz
@@
We start with defining the NJL Lagrangian in eq. 1, within the light quark sector (especially $u$ and $d$ quarks within the exact isospin approximation), with the original aim of describing the low mass mesons such as the pion, the sigma, omega, and the rho mesons:
${\cal L}_{\mathrm{NJL}}$ $=\overline{\psi}\left(i\,\gamma^{\mu}\partial_{\mu}\,-\,m_{q}\right)\,\psi\,+\,\frac{G_{1}}{2}\,\left[\left(\overline{\psi}\psi\right)^{2}\,+\ \left(\overline{\psi}\,i\gamma_{5}\vec{\tau}\,\psi\right)^{2}\right]-\,\frac{G_{2}}{2}\,\left[\left(\overline{\psi}\,\gamma^{\mu}\vec{\tau}\,\psi\right)^{2}\,+\,\left(\overline{\psi}\,\gamma^{\mu}\gamma_{5}\vec{\tau}\,\psi\right)^{2}\,+\,\left(\overline{\psi}\,\gamma^{\mu}\,\psi\right)^{2}\right]~,$ (1)
@@
where $\psi$ are the quark fields. The model depends on three parameters, scalar coupling constants $G_{1}$, vector coupling constant $G_{2}$, and the current quark mass $m_{q}$. In addition to the three parameters from eq. 1, a non-covariant cutoff parameter $\Lambda$ is introduced to avoid UV divergences in integrals, see below. In the absence of $G_{2}$, the parameters ($G_{1},m_{q},$ and $\Lambda$) are adjusted to reproduce the pion mass, the pion decay constant, and the quark condensate. The effect of the meson parameters on $G_{2}$ is discussed later in this section. The choice of the parameters considered in this work is discussed in section III.
@@
The detailed methodology and the procedure of the bosonization of the NJL model used in this work have been developed in Refs. [63, 68]. In summary, the bosonization of the NJL model in the current framework is done using path integral techniques to integrate the quark field, and after a chiral rotation of the quark field, it can be equivalently written in a semi-bosonized form involving a pion field, embedded in the unitary operator $U=\xi^{2}$, a scalar field ($\cal S$}), a vector field ($\Omega_{\mu}$), vector iso-vector ($\vec{V}_{\mu}$), and an axial-vector field ($\vec{A}_{\mu}$[63, 68]. The important virtue of the method lies in the fact that all the parameters entering the bosonized Lagrangian functionally depend on the value of the scalar field $\mathcal{S}$ whose vacuum expectation value coincides with the constituent quark mass $M_{0}\sim 350$ MeV. The amount of partial chiral symmetry restoration is associated with the decrease of the field $\mathcal{S}$. The value of $\mathcal{S}$, or its deviation $s$ from the vacuum value given by $s=(F_{\pi}/M_{0})(\mathcal{S}-M_{0})$, in the dense medium is obtained from the equation of motion for the scalar field as explained in section III.2.2. It follows that all the quantities related to the nucleon properties should be seen as a functional of the nuclear scalar field $s$ induced by the presence of the other nucleons. In uniform matter, the value of the scalar field $\cal S$ is considered as uniform and equal to the value it takes for the density of nuclear matter. The fluctuation of the scalar field in and out of the nucleon is therefore disregarded by this prescription, which is in the spirit of the Born-Oppenheimer approximation in finite-size nuclei [86, 87, 67]. The value of the scalar field for a given nucleon is taken at the nucleon center of mass (CM) position $\bm{R}_{N}$ of the nucleon for the nuclear vector density $\rho(\bm{R}_{N})$. For finite-size nuclei, $\mathcal{S}(\bm{R}_{N})=(M_{0}/F_{\pi})(F_{\pi}+s(\bm{R}_{N}))$.
@@
In the current work, we describe the soliton with the pion, the vector, and the iso-vector mesons. Hence, the effective Lagrangian takes the form (after omitting axial meson contribution and the explicit scalar meson terms of the meson Lagrangian of Ref. [63]):
[…]
@@
where $N_{c}$ is the number of colors and $N_{f}$ is the number of flavors.
Furthermore in eq. 2,
@@
[…]
@@
[…]
In the NJL-soliton framework, we construct baryons as Skyrmions within a chiral model incorporating $\pi$, $\rho$, and $\omega$ meson fields describing the baryon structure. For simplicity, we focus on static solutions, neglecting the (iso-)spin quantization of the soliton. Unlike the standard Skyrme model, which relies on a higher-order Skyrme term for stable soliton solutions, here we focus on the solitons stabilized by the vector $\omega$ and iso-vector $\rho$ mesons [53, 54, 56]. Furthermore, to describe the soliton, we consider the hedgehog ansatz for the pion field, with corresponding configurations for the vector mesons, specifically the Wu-Yang-’t Hooft-Polyakov ansatz for the $\rho$ meson and a time-component-only form for the $\omega$ meson due to the absence of spatial source terms in static solutions:
@@
where $\bm{\tau}$ denotes the Pauli matrices, $F(r)$ is the chiral angle, $\epsilon^{ika}$ is the Levi-Civita symbol with $i$ and $a$ representing spatial and isospin indices, respectively, and $G(r)$ and $\omega(r)$ are the radial profiles of the $\rho$ and $\omega$ mesons. Given the Lagrangian density in eq. 12, and the particular choices of the meson profiles, the soliton energy density ($\varepsilon_{\rm{slt}}$) and pressure ($p_{\rm{slt}}$) can be obtained as outlined in Refs. [56, 57, 54, 88];
$\varepsilon_{\rm{slt}}(r)$ $=F_{\pi}^{2}({\cal S})\left(\frac{1}{2}(F^{\prime})^{2}+\frac{\sin^{2}F}{r^{2}}+\frac{a({\cal S})(G+1-\cos F)^{2}}{r^{2}}\right)+\frac{G^{\prime 2}}{4g_{v}^{2}({\cal S})r^{2}}+\frac{G^{2}(G+2)^{2}}{8g_{v}^{2}({\cal S})r^{4}}$
$\quad+M_{\pi}^{2}({\cal S})F_{\pi}^{2}({\cal S})\ (1-\cos F)-\frac{1}{2}M_{v}^{2}({\cal S})\omega^{2}-\frac{1}{2}(\omega^{\prime})^{2}+\frac{3g_{v}({\cal S})}{2\pi^{2}r^{2}}\omega F^{\prime}\sin^{2}F~,$ (18)
and,
$p_{\rm{slt}}(r)$ $=-\frac{F_{\pi}^{2}({\cal S})}{3r^{2}}\left[\tfrac{1}{2}(rF^{\prime})^{2}+\sin^{2}F+a({\cal S})(G+1-\cos F)^{2}\right]-M_{\pi}^{2}({\cal S})\,F_{\pi}^{2}({\cal S})\,(1-\cos F)$
$+\frac{(G^{\prime})^{2}}{12g_{v}^{2}({\cal S})r^{2}}+\frac{G^{2}(G+2)^{2}}{24g_{v}^{2}({\cal S})r^{4}}+\tfrac{1}{2}M_{v}^{2}({\cal S})\,\omega^{2}-\tfrac{1}{6}(\omega^{\prime})^{2}~.$ (19)
In eqs. 18 and 19, $F_{\pi}({\cal S})$, $M_{\pi}({\cal S})$, $M_{v}({\cal S})$, $g_{v}({\cal S})$, and $a({\cal S})$ are related to the NJL model parameters and the scalar field $S$ as described in eqs. (8), (7), (10), (5), and (11), respectively. For clarity and simplicity, we have omitted the subscript $\rm NJL$ afterwards when representing the meson parameters, similar to those in eqs. 18 and 19.
@@
Hence, in the current framework, the effects of chiral symmetry restoration at high density enter the soliton description solely through the scalar field $\cal S$}. Throughout this article, $f_{\pi}\equiv F_{\pi,\rm NJL}({\cal S}=M_{0})$ represents the pion decay constant parameter in the vacuum if not mentioned elsewhere. The notation is chosen to be compatible with the general notation.
II.2 Field Equations and Numerical Method
@@
[…]
@@
with the quantized baryon charge $B=\int_{0}^{\infty}4\pi r^{2}\rho_{B}(r)\,dr$. For $B=1$ solutions and ensuring finite energy of the soliton eq. 18, the boundary conditions are:
@@
The proportionality constant in eq. 30, does not depend upon $r$ and is adjusted to normalize the iso-scalar density ($\int\ \rho_{S}(r)\ d^{3}r=1$). To extract nucleon properties, one must solve the coupled second-order differential equations given in eqs. 23, 24 and 25) subject to boundary conditions at the origin and at spatial infinity. The known initial values areF(0)F(0),$G(0)$, and $\omega^{\prime}(0)$, while the derivatives $F^{\prime}(0)$, $G^{\prime}(0)$, and the initial value $\omega(0)$ must be determined iteratively. A common strategy is to use the shooting method: integrating outward from $r=0$ with trial values for the unknowns, adjusting them until the boundary conditions at large $r$ are satisfied.
@@
In practice, this system is highly stiff and extremely sensitive to initial conditions. Even a small deviations in the trial parameters lead to divergent or oscillatory (“frog-like”) behavior, hindering the ability to reach physically meaningful solutions at $r\gtrsim 3\,\mathrm{fm}$. Although implicit methods offer better control over instabilities compared to explicit methods, such as the Runge–Kutta (RK4) method, they still fail to prevent exponential growth in the solutions. Moreover, even when a stable profile is achieved up to a finite matching point $r_{\infty}$ with suitable values of $F^{\prime}(0)$, $G^{\prime}(0)$, and $\omega(0)$, extending the integration often results in divergent behavior. While we can reproduce the soliton solutions presented in Ref. [58] using the shooting method with fine-tuned initial conditions, the system’s extreme sensitivity renders the shooting method impractical for systematic studies.
@@
The system of coupled, second-order differential equations for the profile functions $F(r)$, $G(r)$, and $\omega(r)$ constitutes a mixture of initial value problem and boundary value problem (somewhat similar in structure to those encountered in quantum mechanics, such as the inverse-cusp potential). Due to the stiffness and sensitivity of the equations, conventional shooting methods prove inadequate. Instead, we employ a relaxation method from Ref. [89], which is naturally suited for solving such problems.
II.3 Numerical Implementations

### v2

@@ 2 Theoretical Framework
To construct the in-medium modified nucleon state and quark core state, we follow an approach similar to that adopted in the NJL version of the chiral confining model (CCM), as briefly sketched below. The chiral confining–NJL model captures the essential dynamics of chiral symmetry breaking and quark confinement, providing an effective framework to describe low-energy QCD phenomena. It elucidates the response of composite nucleons to the nuclear scalar field and the interplay between confinement and chiral symmetry breaking in the in-medium evolution of the nucleon mass. However, in the present work, we restrict our analysis to the NJL-soliton framework for describing the nucleon, without introducing an explicit confining force such as that associated with a string tension. Instead, the role of confinement is effectively mimicked by mesons (with the soliton stabilized by vector mesons in the current framework) within the soliton picture. A description of the nucleon, including an explicit confining mechanism, will be addressed in future work. As detailed in Ref. Chanfray2024EPJA ; Chanfray2025 , the CCM provides a robust theoretical framework for understanding QCD phenomenology, particularly in bridging QCD with the description of nuclear and NS matter. In its NJL version, this model builds upon the NJL model, whose bosonization procedure and theoretical formulation are comprehensively presented in Ref. Chanfray_PRC2011 ; Chanfray2025 . We adopt the same approach here.
2.1 Bosonized NJL Model and Chiral Soliton Ansatz
@@
We start with defining the NJL Lagrangian in Eq. (1), within the light quark sector (especially $u$ and $d$ quarks within the exact isospin approximation), with the original aim of describing the low mass mesons such as the pion, the sigma, omega, and the rho mesons:
${\cal L}_{\mathrm{NJL}}$ $=\overline{\psi}\left(i\,\gamma^{\mu}\partial_{\mu}\,-\,m_{q}\right)\,\psi\,$
$+\,\frac{G_{1}}{2}\,\left[\left(\overline{\psi}\psi\right)^{2}\,+\ \left(\overline{\psi}\,i\gamma_{5}\vec{\tau}\,\psi\right)^{2}\right]$
$-\,\frac{G_{2}}{2}\,\left[\left(\overline{\psi}\,\gamma^{\mu}\vec{\tau}\,\psi\right)^{2}\,+\,\left(\overline{\psi}\,\gamma^{\mu}\gamma_{5}\vec{\tau}\,\psi\right)^{2}\,+\,\left(\overline{\psi}\,\gamma^{\mu}\,\psi\right)^{2}\right]~,$ (1)
@@
where $\psi$ are the quark fields. The model depends on three parameters, scalar coupling constants $G_{1}$, vector coupling constant $G_{2}$, and the current quark mass $m_{q}$. In addition to the three parameters from Eq. (1), a non-covariant cutoff parameter $\Lambda$ is introduced to avoid UV divergences in integrals, see below. In the absence of $G_{2}$, the parameters ($G_{1},m_{q},$ and $\Lambda$) are adjusted to reproduce the pion mass, the pion decay constant, and the quark condensate. The effect of the meson parameters on $G_{2}$ is discussed later in this section. The choice of the parameters considered in this work is discussed in Section 3.
@@
The detailed methodology and the procedure of the bosonization of the NJL model used in this work have been developed in Refs. Chanfray_PRC2011 ; Chanfray2025 . In summary, the bosonization of the NJL model in the current framework is done using path integral techniques to integrate the quark field, and after a chiral rotation of the quark field, it can be equivalently written in a semi-bosonized form involving a pion field, embedded in the unitary operator $U=\xi^{2}$, a scalar field ($\cal S$}), a vector field ($\Omega_{\mu}$), vector iso-vector ($\vec{V}_{\mu}$), and an axial-vector field ($\vec{A}_{\mu}$Chanfray_PRC2011 ; Chanfray2025 . The important virtue of the method lies in the fact that all the parameters entering the bosonized Lagrangian functionally depend on the value of the scalar field $\mathcal{S}$ whose vacuum expectation value coincides with the constituent quark mass $M_{0}\sim 350$ MeV. The amount of partial chiral symmetry restoration is associated with the decrease of the field $\mathcal{S}$. The value of $\mathcal{S}$, or its deviation $s$ from the vacuum value given by $s=(F_{\pi}/M_{0})(\mathcal{S}-M_{0})$, in the dense medium is obtained from the equation of motion for the scalar field as explained inSection 3.2.2. It follows that all the quantities related to the nucleon properties should be seen as a functional of the nuclear scalar field $s$ induced by the presence of the other nucleons. In uniform matter, the value of the scalar field $\cal S$ is considered as uniform and equal to the value it takes for the density of nuclear matter. The fluctuation of the scalar field in and out of the nucleon is therefore disregarded by this prescription, which is in the spirit of the Born-Oppenheimer approximation in finite-size nuclei GuichonNuPhA ; GuichonPrPNP ; Chanfray2024EPJA . The value of the scalar field for a given nucleon is taken at the nucleon center of mass (CM) position $\bm{R}_{N}$ of the nucleon for the nuclear vector density $\rho(\bm{R}_{N})$. For finite-size nuclei, $\mathcal{S}(\bm{R}_{N})=(M_{0}/F_{\pi})(F_{\pi}+s(\bm{R}_{N}))$.
@@
In the current work, we describe the soliton with the pion, the vector, and the iso-vector mesons. Hence, the effective Lagrangian takes the form (after omitting axial meson contribution and the explicit scalar meson terms of the meson Lagrangian of Ref. Chanfray_PRC2011 ):
${\cal L}_{\mathrm{mes}}$ $=\,\frac{m_{q}\,{\cal S}}{4\,G_{1}}\,\mathrm{Tr}_{f}(U\,+\,U^{\dagger}\,-\,2)\,$
$+\,\tilde{I}({\cal S})\,{\cal S}^{2}\,\mathrm{Tr}_{f}\left({\cal A}^{\mu}_{c}\cdot{\cal A}^{c}_{\mu}\right)$
$+\,\frac{1}{4\,G_{2}}\,\mathrm{Tr}_{f}\left({V}^{\mu}+{\cal V}^{\mu}_{c}\right)^{2}\$
@@
where $N_{c}$ is the number of colors and $N_{f}$ is the number of flavors. Furthermore in Eq. (2),
@@
${\cal L}_{\rm slt}$ $=\,\frac{1}{4}F^{2}_{\pi}({\cal S})M^{2}_{\pi}({\cal S})\,\ \mathrm{Tr}_{f}(U\,+\,U^{\dagger}\,-\,2)\,$
$+\,\frac{F_{\pi}^{2}({\cal S})}{4}\mathrm{Tr}_{f}\left(\partial_{\mu}U\partial^{\mu}U^{\dagger}\right)\,$
$+\,a({\cal S})F_{\pi}^{2}({\cal S})\,\mathrm{Tr}_{f}\left({V}^{\mu}+{\cal V}^{\mu}_{c}\right)^{2}$
@@
$B^{\mu}=\frac{1}{24\pi^{2}}\epsilon^{\mu\nu\alpha\beta}\ \mathrm{Tr}\left[U^{\dagger}\partial_{\nu}U\ .\ U^{\dagger}\partial_{\alpha}U\ .\ U^{\dagger}\partial_{\beta}U\right]~.$ (14)
The solitons considered in this work are constructed within a bosonized NJL framework using a derivative expansion, whose domain of validity is well understood Zuckert1995b . Although this approximation is not systematically controlled when the soliton size approaches the scale of integrated-out physics, it provides a standard and widely used description of the leading mesonic dynamics responsible for soliton formation. While more refined treatments have been developed to address this limitation (see for instance  Diakonov1997a ; Diakonov1997b ; Goeke2001 ), their implementation lies beyond the scope of the present work. However, our aim here is not to refine or study the detailed chiral soliton model itself, but to illustrate a physical mechanism within a simple and tractable framework (a framework that has been adopted in  Fukushima2020 ) that retains the essential chiral symmetry properties.
In the NJL-soliton framework, we construct baryons as Skyrmions within a chiral model incorporating $\pi$, $\rho$, and $\omega$ meson fields describing the baryon structure. For simplicity, we focus on static solutions, neglecting the (iso-)spin quantization of the soliton. Unlike the standard Skyrme model, which relies on a higher-order Skyrme term for stable soliton solutions, here we focus on the solitons stabilized by the vector $\omega$ and iso-vector $\rho$ mesons ADKINS1984251 ; IGARASHI1985 ; Meissner1986 . Furthermore, to describe the soliton, we consider the hedgehog ansatz for the pion field, with corresponding configurations for the vector mesons, specifically the Wu-Yang-’t Hooft-Polyakov ansatz for the $\rho$ meson and a time-component-only form for the $\omega$ meson due to the absence of spatial source terms in static solutions:
@@
where $\bm{\tau}$ denotes the Pauli matrices, $F(r)$ is the chiral angle, $\epsilon^{ika}$ is the Levi-Civita symbol with $i$ and $a$ representing spatial and isospin indices, respectively, and $G(r)$ and $\omega(r)$ are the radial profiles of the $\rho$ and $\omega$ mesons. Given the Lagrangian density in Eq. (12), and the particular choices of the meson profiles, the soliton energy density ($\varepsilon_{\rm{slt}}$) can be obtained as outlined in Refs. Meissner1986 ; Meissner1987 ; IGARASHI1985 ; Ma2013 ;
$\varepsilon_{\rm{slt}}(r)$ $=F_{\pi}^{2}({\cal S})\left(\frac{1}{2}(F^{\prime})^{2}+\frac{\sin^{2}F}{r^{2}}\right)$
$+F_{\pi}^{2}({\cal S})\left(\frac{a({\cal S})(G+1-\cos F)^{2}}{r^{2}}\right)$
$+\frac{G^{\prime 2}}{4g_{v}^{2}({\cal S})r^{2}}+\frac{G^{2}(G+2)^{2}}{8g_{v}^{2}({\cal S})r^{4}}$
$-\frac{1}{2}M_{v}^{2}({\cal S})\omega^{2}-\frac{1}{2}(\omega^{\prime})^{2}+\frac{3g_{v}({\cal S})}{2\pi^{2}r^{2}}\omega F^{\prime}\sin^{2}F$~
$+M_{\pi}^{2}({\cal S})F_{\pi}^{2}({\cal S})\ (1-\cos F)~.$ (18)
As pointed out in Ref. Fukushima2025 , there is a pseudo-gauge ambiguity impacting the nucleon EOS. However, in this work, we follow the pragmatic prescription of  Fukushima2020 by employing a mixture of the canonical and Belinfante energy momentum tensor (EMT), which removes the singularity in the pressure (see Ref. Fukushima2025 for details). Within this framework and consideration, the soliton pressure ($p_{\rm slt}$) is calculated as in  Fukushima2020 :
$p_{\rm{slt}}(r)$ $=-\frac{F_{\pi}^{2}({\cal S})}{3r^{2}}\left[\frac{1}{2}(rF^{\prime})^{2}+\sin^{2}F\right]$
$-\frac{F_{\pi}^{2}({\cal S})}{3r^{2}}\left[a({\cal S})(G+1-\cos F)^{2}\right]$
$+\frac{(G^{\prime})^{2}}{12g_{v}^{2}({\cal S})r^{2}}+\frac{G^{2}(G+2)^{2}}{24g_{v}^{2}({\cal S})r^{4}}$
$+\tfrac{1}{2}M_{v}^{2}({\cal S})\,\omega^{2}-\tfrac{1}{6}(\omega^{\prime})^{2}$
$-M_{\pi}^{2}({\cal S})\,F_{\pi}^{2}({\cal S})\,(1-\cos F)~.$ (19)
In Eqs. (18) and  (19), $F_{\pi}({\cal S})$, $M_{\pi}({\cal S})$, $M_{v}({\cal S})$, $g_{v}({\cal S})$, and $a({\cal S})$ are related to the NJL model parameters and the scalar field $S$ as described in Eqs. (8), (7), (10), (5), and (11), respectively. For clarity and simplicity, we have omitted the subscript $\rm NJL$ afterwards when representing the meson parameters, similar to those in Eqs. (18) and  (19).
@@
The scalar field $\cal S$} is introduced as a chiral-invariant effective field that parametrizes medium-induced modifications of the underlying chiral dynamics. It is worth mentioning that the scalar field ($\cal S$) is not an order parameter of chiral symmetry breaking, unlike the quark condensate, whose density dependence arises from both the scalar mean field and the pionic scalar density. Nevertheless, $\cal S$} remains a useful and physically meaningful indicator of the in-medium evolution of chiral dynamics, as it is closely related to the constituent quark mass in NJL-based approaches. Hence, in the present framework, medium effects associated with partial chiral symmetry restoration enter the soliton description through the density dependence of $\cal S$. As the density increases and chiral symmetry is progressively restored (corresponding to $|s|\to f_{\pi}$ or ${\cal S}\to 0$ ), the in-medium pion and scalar masses approach each other ( see discussions in Section 3.2.2 and  C). This signals the gradual degeneracy of chiral partners and reflects the progressive restoration of chiral symmetry. This behavior is consistent with general expectations from chiral symmetry restoration and provides an interpretation of the role played by $\cal S$ in the model.
Throughout this article, $f_{\pi}\equiv F_{\pi,\rm NJL}({\cal S}=M_{0})$ represents the pion decay constant parameter in the vacuum if not mentioned elsewhere. The notation is chosen to be compatible with the general notation.
2.2 Field Equations and Numerical Method
@@
$F^{\prime\prime}$ $=-\frac{2F^{\prime}}{r}+\frac{\sin{2F}}{r^{2}}[1-{a({\cal S})}]+\frac{2{a({\cal S})}\sin F}{r^{2}}(G+1)$~
@@
with the quantized baryon charge $B=\int_{0}^{\infty}4\pi r^{2}\rho_{B}(r)\,dr$. For $B=1$ solutions and ensuring finite energy of the soliton Eq. (18), the boundary conditions are:
@@
The proportionality constant in Eq. (30), does not depend upon $r$ and is adjusted to normalize the iso-scalar density ($\int\ \rho_{S}(r)\ d^{3}r=1$). To extract nucleon properties, one must solve the coupled second-order differential equations given in Eqs. (23), (24), and  (25).) subject to boundary conditions at the origin and at spatial infinity. The known initial values areF(0)F(0),$G(0)$, and $\omega^{\prime}(0)$, while the derivatives $F^{\prime}(0)$, $G^{\prime}(0)$, and the initial value $\omega(0)$ must be determined iteratively. A common strategy is to use the shooting method: integrating outward from $r=0$ with trial values for the unknowns, adjusting them until the boundary conditions at large $r$ are satisfied.
@@
In practice, this system is highly stiff and extremely sensitive to initial conditions. Even a small deviations in the trial parameters lead to divergent or oscillatory (“frog-like”) behavior, hindering the ability to reach physically meaningful solutions at $r\gtrsim 3\,\mathrm{fm}$. Although implicit methods offer better control over instabilities compared to explicit methods, such as the Runge–Kutta (RK4) method, they still fail to prevent exponential growth in the solutions. Moreover, even when a stable profile is achieved up to a finite matching point $r_{\infty}$ with suitable values of $F^{\prime}(0)$, $G^{\prime}(0)$, and $\omega(0)$, extending the integration often results in divergent behavior. While we can reproduce the soliton solutions presented in Ref. Fukushima2020 using the shooting method with fine-tuned initial conditions, the system’s extreme sensitivity renders the shooting method impractical for systematic studies.
@@
The system of coupled, second-order differential equations for the profile functions $F(r)$, $G(r)$, and $\omega(r)$ constitutes a mixture of initial value problem and boundary value problem (somewhat similar in structure to those encountered in quantum mechanics, such as the inverse-cusp potential). Due to the stiffness and sensitivity of the equations, conventional shooting methods prove inadequate. Instead, we employ a relaxation method from Ref. numerical , which is naturally suited for solving such problems.
2.3 Numerical Implementations